OpenStudy (anonymous):

The first derivative of the function f is given by f(x) = (cos^2(x))/(x) - (1/5). How many critical values does f have in the open interval (0, 10)?

OpenStudy (tkhunny):

You did not clarify your last post. Which is it and what's your plan?

OpenStudy (anonymous):

I don't know what you're asking.

OpenStudy (anonymous):

What last post?

OpenStudy (tkhunny):

Is it \(\dfrac{\cos^{2}(x)}{x}-\dfrac{1}{5}\;or\;\dfrac{\cos^{2}(x)}{x-\tiny{\dfrac{1}{5}}}\)?